# CarPlanner: Consistent Auto-regressive Trajectory Planning for Large-scale Reinforcement Learning in Autonomous Driving

### v1

@@ Abstract
footnotetext: Project lead. ${}^{\textrm{\Letter}}$ Corresponding author (rxiong@zju.edu.cn).
@@ 1 Introduction
The first category, illustrated in Fig. 1 (a), involves generating an initial trajectory estimate and subsequently refining it through iterative applications of RL. However, recent studies, including Gen-Drive [21], suggest that it continues to lag behind SOTA IL and rule-based planners. One notable limitation of this approach is its neglect of the temporal causality inherent in the trajectory planning task. Additionally, the complexity of direct optimization over high-dimensional trajectory space can hinder the performance of RL algorithms. The second category consists of auto-regressive models, shown in Fig. 1 (b), which generate the poses of the ego vehicle recurrently using a single-step policy within a world model. In this category, ego poses at all time steps are consolidated to form the overall planned trajectory. As taking temporal causality into account, current auto-regressive models allow for interactive behaviors. However, a common limitation is their reliance on auto-regressively random sampling from action distributions to generate multi-modal trajectories. This vanilla auto-regressive procedure may compromise long-term consistency and unnecessarily expand the exploration space in RL, leading to poor performance.
@@ 2 Related Work
Figure 2: CarPlanner contains four parts. (1) The non-reactive world model takes initial state $\boldsymbol{s}_{0}$ as input and predicts the future trajectories of traffic agents. (2) The mode selector outputs scores based on the initial state and the modes $\boldsymbol{c}$. (3) The trajectory generator obeys an auto-regressive structure condition on the consistent mode and produces mode-aligned multi-modal trajectories. (4) The rule-augmented selector compensates the mode scores by safety, comfort, and progress metrics.
@@ 3.2 Problem Formulation
The state sequence can be further formulated in an auto-regressive fashion and decomposed into policy and world model:
$\begin{split}&P(\boldsymbol{s}_{0},\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{T})=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(s^{0}_{t+1},s^{1:N}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}\underbrace{\pi(a_{t}|\boldsymbol{s}_{t})}_{\text{Policy}}\underbrace{P(s^{1:N}_{t+1}|\boldsymbol{s}_{t})}_{\text{World Model}}.\\[-4.26773pt] \end{split}$ (3)
@@
Since we focus on the ego trajectory planning, the consistent mode $\boldsymbol{c}$ does not impact world model.
@@
Non-reactive world model. The world model formulated in LABEL:equation:car needs to be employed in every time step since it produces the poses of traffic agents at time step $t+1$ based on current state $\boldsymbol{s}_{t}$. In practice, this process is time-consuming and we do not observe a performance improvement by using this world model, therefore, we use trajectory predictors $P(s^{1:N}_{1:T}|\boldsymbol{s}_{0})$ as non-reactive world model that produces all future poses of traffic agents in one shot given initial state $\boldsymbol{s}_{0}$.
@@ 3.3 Planner Architecture
The framework of our proposed CarPlanner is illustrated in Fig. 2, comprising four key components: 1) the non-reactive world model, 2) the mode selector, 3) the trajectory generator, and 4) the rule-augmented selector. Our planner operates within a generation-selection framework. Given an initial state $s_{0}$ and all possible $N_{\text{mode}}$ modes, the trajectory selector evaluates and assigns scores to each mode. The trajectory generator then produces $N_{\text{mode}}$ trajectories that correspond to their respective modes.
@@
3.3.1 Non-reactive World Model
@@ 3.3.3 Trajectory Generator
Invariant-view module (IVM). Before feeding the mode and state into the network, we preprocess them to eliminate time information. For the map and agent information in state $\boldsymbol{s}_{t}$, we select the $K$ nearest neighbors (KNN) [28] to the ego current pose and only feed these into the policy. $K$ is set to the half of map and agent elements respectively. Regarding the routes that capture lateral behaviors, we filter out the segments where the point closest to the current pose of the ego vehicle is the starting point, retaining $K_{r}$ points. In this case, $K_{r}$ is set to a quarter of $N_{r}$ points in one route. Finally, we transform the routes, agent, and map poses into the coordinate frame of the ego vehicle at the current time step $t$. We then subtract the historical time steps $t-H:t$ from the current time step $t$, yielding time steps in the range $-H:0$.
@@ 3.4 Training
We first train the non-reactive world model and freeze the weights during the training of the mode selector and trajectory generator. Instead of feeding all modes to the generator, we apply a winner-takes-all strategy, wherein a positive mode is assigned based on the ego ground-truth trajectory and serves as a condition for the trajectory generator.
@@
Loss function. For training the selector, we use cross-entropy loss that is the negative log-likelihood of the positive mode. The PPO loss consists of policy improvement loss, value estimation loss, and entropy loss for exploration. The full description can be found in the supplementary.
@@
Ego-history dropout. Previous works [1, 13, 5, 4] suggest that planners trained via IL may rely too heavily on past poses and neglect environmental state information. To counter this, we combine techniques from ChauffeurNet [1] and PlanTF [5] into an ego-history dropout module, randomly masking velocity and acceleration during IL training, but not during RL training. Further discussions can be found in Sec. 4.4.
@@ 4.1 Experimental Setup
Design Choices Closed-loop metrics ($\uparrow$) Open-loop metrics ($\downarrow$)
@@
Table 3: Ablation studies on the design choices in RL training. Experimental results are based on the Test14-random non-reactive benchmark.
We use the closed-loop score (CLS) provided by the official nuPlan devkithttps://github.com/motional/nuplan-devkit to assess the performance of all methods. The CLS score comprehends different aspects such as safety (S-CR, S-TTC), drivable area compliance (S-Area), progress (S-PR), comfort, etc. Based on the different behavior types of traffic agents, CLS is detailed into CLS-NR (non-reactive) and CLS-R (reactive). To further analyze various components of our methods, we also use open-loop metrics, such as loss of trajectory generator and selector.
4.2 Comparison with SOTAs
@@ 4.3 Ablation Studies
Design Choices Closed-loop metrics ($\uparrow$) Open-loop metrics ($\downarrow$)
Loss Type Mode Dropout Scorer Side Task Ego-history Dropout Backbone Sharing CLS-NR S-CR S-Area S-PR S-Comfort Loss Selector Loss Generator
@@
Table 4: Effect of different components on IL and RL Loss using our CarPlanner. Experimental results are based on the Test14-random non-reactive benchmark.
@@
Influence of reward items. The results demonstrate that the DE and quality rewards are complementary. When using the DE reward only, the planner tends to generate static trajectories and achieves a low progress metric. This occurs because the ego vehicle begins in a safe, drivable state, but moving forward is at risk of collisions or leaving the drivable area. On the other hand, compared to using DE reward only, incorporating the quality reward significantly improves closed-loop metrics. For instance, the S-CR metric rises from 97.49 to 99.22, and the S-Area metric rises from 96.91 to 99.22. These improvements indicate that the quality reward encourages safe and comfortable behaviors.
@@ 4.4 Extention to IL
In addition to designing for RL training, we also extend the CarPlanner to incorporate IL. We conduct rigorous analysis to compare the effects of various design choices in IL and RL training, as summarized in Tab. 4. Our findings indicate that while mode dropout and scorer side task contribute to both IL and RL training, ego-history dropout and backbone sharing, often effective in IL, are less suitable for RL.
Ego-history dropout. This technique effectively addresses the causal confusion issue in IL by preventing over-reliance on ego history, leading to more generalized planning behavior. Our experiments confirm that ego-history dropout positively impacts IL training, as it improves performance across closed-loop metrics like S-CR and S-Area. However, in RL training, we observe a negative impact on advantage estimation due to ego-history dropout, which significantly affects the accuracy of value estimation, leading to closed-loop performance degradation. This indicates that RL training naturally overcomes the causal confusion issue of IL, since explicit task-oriented metrics can be incorporated into reward functions, highlighting the potential of RL to pushing the boundaries of learning-based planning.
@@ 4.5 Qualitative Results
We provide qualitative results as shown in Fig. 3. In this scenario, ego vehicle is required to execute a right turn while navigating around pedestrians. In this case, Our method demonstrates a smooth, efficient performance. From $t_{\text{sim}}=0s$ to $t_{\text{sim}}=9s$, all methods wait for the pedestrians to cross the road. At $t_{\text{sim}}=10s$, an unexpected pedestrian goes back and prepares to re-cross the road. PDM-Closed is unaware of this situation and takes an emergency stop but still intersects with this pedestrian. In contrast, our IL variant displays an awareness of the pedestrian’s movements and consequently conducts a braking maneuver. However, it still remains close to the pedestrian. Our RL method avoids this hazard by starting up early up to $t_{\text{sim}}=9s$ and achieves the highest progress and safety metrics.

### v2

@@ Abstract
footnotetext: Project lead. ${}^{\textrm{\Letter}}$ Corresponding author.
@@ 1 Introduction
The first category, illustrated in Fig. 1 (a), involves generating an initial trajectory estimate and subsequently refining it through iterative applications of RL. However, recent studies, including Gen-Drive [21], suggest that it continues to lag behind SOTA IL and rule-based planners. One notable limitation of this approach is its neglect of the temporal causality inherent in the trajectory planning task. Additionally, the complexity of direct optimization over high-dimensional trajectory space can hinder the performance of RL algorithms. The second category consists of auto-regressive models, shown in Fig. 1 (b), which generate the poses of the ego vehicle recurrently using a single-step policy within a transition model. In this category, ego poses at all time steps are consolidated to form the overall planned trajectory. As taking temporal causality into account, current auto-regressive models allow for interactive behaviors. However, a common limitation is their reliance on auto-regressively random sampling from action distributions to generate multi-modal trajectories. This vanilla auto-regressive procedure may compromise long-term consistency and unnecessarily expand the exploration space in RL, leading to poor performance.
@@ 2 Related Work
Figure 2: CarPlanner contains four parts. (1) The non-reactive transition model takes initial state $\boldsymbol{s}_{0}$ as input and predicts the future trajectories of traffic agents. (2) The mode selector outputs scores based on the initial state and the modes $\boldsymbol{c}$. (3) The trajectory generator obeys an auto-regressive structure condition on the consistent mode and produces mode-aligned multi-modal trajectories. (4) The rule-augmented selector compensates the mode scores by safety, comfort, and progress metrics.
@@ 3.2 Problem Formulation
The state sequence can be further formulated in an auto-regressive fashion and decomposed into policy and transition model:
$\begin{split}&P(\boldsymbol{s}_{0},\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{T})=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(\boldsymbol{s}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}P(s^{0}_{t+1},s^{1:N}_{t+1}|\boldsymbol{s}_{t})\\ &=\rho_{0}(\boldsymbol{s}_{0})\prod_{t=0}^{T-1}\underbrace{\pi(a_{t}|\boldsymbol{s}_{t})}_{\text{Policy}}\underbrace{P(s^{1:N}_{t+1}|\boldsymbol{s}_{t})}_{\text{Transition Model}}.\\[-4.26773pt] \end{split}$ (3)
@@
Since we focus on the ego trajectory planning, the consistent mode $\boldsymbol{c}$ does not impact transition model.
@@
Non-reactive transition model. The transition model formulated in LABEL:equation:car needs to be employed in every time step since it produces the poses of traffic agents at time step $t+1$ based on current state $\boldsymbol{s}_{t}$. In practice, this process is time-consuming and we do not observe a performance improvement by using this transition model, therefore, we use trajectory predictors $P(s^{1:N}_{1:T}|\boldsymbol{s}_{0})$ as non-reactive transition model that produces all future poses of traffic agents in one shot given initial state $\boldsymbol{s}_{0}$.
@@ 3.3 Planner Architecture
The framework of our proposed CarPlanner is illustrated in Fig. 2, comprising four key components: 1) the non-reactive transition model, 2) the mode selector, 3) the trajectory generator, and 4) the rule-augmented selector. Our planner operates within a generation-selection framework. Given an initial state $s_{0}$ and all possible $N_{\text{mode}}$ modes, the trajectory selector evaluates and assigns scores to each mode. The trajectory generator then produces $N_{\text{mode}}$ trajectories that correspond to their respective modes.
@@
3.3.1 Non-reactive Transition Model
@@ 3.3.3 Trajectory Generator
Invariant-view module (IVM). Before feeding the mode and state into the network, we preprocess them to eliminate time information. For the map and agent information in state $\boldsymbol{s}_{t}$, we select the $K$ nearest neighbors (KNN) [29] to the ego current pose and only feed these into the policy. $K$ is set to the half of map and agent elements respectively. Regarding the routes that capture lateral behaviors, we filter out the segments where the point closest to the current pose of the ego vehicle is the starting point, retaining $K_{r}$ points. In this case, $K_{r}$ is set to a quarter of $N_{r}$ points in one route. Finally, we transform the routes, agent, and map poses into the coordinate frame of the ego vehicle at the current time step $t$. We subtract the historical time steps $t-H:t$ from the current time step $t$, yielding time steps in range $-H:0$.
@@ 3.4 Training
We first train the non-reactive transition model and freeze the weights during the training of the mode selector and trajectory generator. Instead of feeding all modes to the generator, we apply a winner-takes-all strategy, wherein a positive mode is assigned based on the ego ground-truth trajectory and serves as a condition for the trajectory generator.
@@
Loss function. For the selector, we use cross-entropy loss that is the negative log-likelihood of the positive mode and a side task that regresses the ego ground-truth trajectory. For the generator, we use PPO [36] loss that consists of three parts: policy improvement, value estimation, and entropy. The full description can be found in the supplementary.
@@ 4.1 Experimental Setup
We use the closed-loop score (CLS) provided by the official nuPlan devkithttps://github.com/motional/nuplan-devkit to assess the performance of all methods. The CLS score comprehends different aspects such as safety (S-CR, S-TTC), drivable area compliance (S-Area), progress (S-PR), comfort, etc. Based on the different behavior types of traffic agents, CLS is detailed into CLS-NR (non-reactive) and CLS-R (reactive).
@@
4.2 Comparison with SOTAs
Design Choices Closed-loop metrics ($\uparrow$) Open-loop losses ($\downarrow$)
@@ 4.2 Comparison with SOTAs
Table 3: Ablation studies on the design choices in RL training. Results are in Test14-random non-reactive benchmark.
@@ 4.3 Ablation Studies
Design Choices Closed-loop metrics ($\uparrow$) Open-loop losses ($\downarrow$)
Loss Type Mode Dropout Selector Side Task Ego-history Dropout Backbone Sharing CLS-NR S-CR S-Area S-PR S-Comfort Loss Selector Loss Generator
@@
Table 4: Effect of different components on IL and RL loss using our CarPlanner. Results are in Test14-random non-reactive benchmark.
@@
Influence of reward items. When using the quality reward only, the planner tends to generate static trajectories and achieves a low progress metric. This occurs because the ego vehicle begins in a safe, drivable state, but moving forward is at risk of collisions or leaving the drivable area. On the other hand, when the quality reward is incorporated alongside the DE reward, it leads to significant improvements in closed-loop metrics compared to using the DE reward alone. For instance, the S-CR metric rises from 97.49 to 99.22, and the S-Area metric rises from 96.91 to 99.22. These improvements indicate that the quality reward encourages safe and comfortable behaviors.
@@ 4.4 Extention to IL
In addition to designing for RL training, we also extend the CarPlanner to incorporate IL. We conduct rigorous analysis to compare the effects of various design choices in IL and RL training, as summarized in Tab. 4. Our findings indicate that while mode dropout and selector side task contribute to both IL and RL training, ego-history dropout and backbone sharing, often effective in IL, are less suitable for RL.
Ego-history dropout. Previous works [1, 13, 5, 4] suggest that planners trained via IL may rely too heavily on past poses and neglect environmental state information. To counter this, we combine techniques from ChauffeurNet [1] and PlanTF [5] into an ego-history dropout module, randomly masking ego history poses and current velocity to alleviate the causal confusion issue.
Our experiments confirm that ego-history dropout benifits IL training, as it improves performance across closed-loop metrics like S-CR and S-Area. However, in RL training, we observe a negative impact on advantage estimation due to ego-history dropout, which significantly affects the value part of generator loss, leading to closed-loop performance degradation. This suggests that RL training naturally addresses the causal confusion problem inherent in IL by uncovering causal relationships that align with the reward signal, which explicitly encodes task-oriented preferences. This capability highlights the potential of RL to push the boundaries of learning-based planning.
@@ 4.5 Qualitative Results
We provide qualitative results as shown in Fig. 3. In this scenario, ego vehicle is required to execute a right turn while navigating around pedestrians. In this case, Our method shows a smooth, efficient performance. From $t_{\text{sim}}=0s$ to $t_{\text{sim}}=9s$, all methods wait for the pedestrians to cross the road. At $t_{\text{sim}}=10s$, an unexpected pedestrian goes back and prepares to re-cross the road. PDM-Closed is unaware of this situation and takes an emergency stop but still intersects with this pedestrian. In contrast, our IL variant displays an awareness of the pedestrian’s movements and consequently conducts a braking maneuver. However, it still remains close to the pedestrian. Our RL method avoids this hazard by starting up early up to $t_{\text{sim}}=9s$ and achieves the highest progress and safety metrics.